Doxycyline oral suspension has the following reconstitution instructions: Add 47.6 mL of water to make a total volume of 60mL. When reconstituted as directed, each teaspoonful (5mL) contains doxycycline monohydrate equivalent to 25mg of doxycycline. The order is for 35mg p.o. b.i.d. for the 9 year old child who weighs 75lbs. How many mL will the nurse administer to the child per dose?
Answer #1

Doxycycline oral suspension comes in powder form. This should be diluted with 47.6mL of distilled water to make a solution of 60 mL.

5mL contains 25mg of doxycycline.

Therefore,

1ml = 25 /5 =5mg of doxycycline.

Ordered dose of doxycycline is 35mg PO b.i.d

Ordered dose in mL = 35 / 5= 7mL

The nurse has to administer 7mL of doxycycline syrup to the child for per dose.